OpenStudy (anonymous):

f(x)=2x^2; a. find f inverse c. evaluate df/dx at x=5 and df(inverse)/dx at x=f(5)

OpenStudy (lgbasallote):

f(x) = 2x^2 to find the inverse..the first step is to change f(x) into y, yes?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

so you ahve y = 2x^2 now...solve for x..do you know how to?

OpenStudy (anonymous):

\[y=\sqrt{x/2}\]

OpenStudy (lgbasallote):

what did you do? why did you swap the variables?

OpenStudy (lgbasallote):

remember... if you start out with y = then the inverse would be x = got it?

OpenStudy (anonymous):

\[f ^{-1}=\sqrt{x/2}\]

OpenStudy (anonymous):

\[f ^{-1}(x)=\sqrt{x/2}\]

OpenStudy (lgbasallote):

nevermind what i said earlier...my source (ahem @TAKEBACKMATH ) told me the inverse should be x =... \[y = \sqrt{\frac x2}\] ^that's right

OpenStudy (anonymous):

that's not really the part I needed help with...it is evaluating the derivative at x=5 and the inverse derivative at x=f(5)....do you know how to do that?

OpenStudy (lgbasallote):

do you know the derivative of 2x^2?

OpenStudy (anonymous):

4x?

OpenStudy (lgbasallote):

right. now substitute x = 5

OpenStudy (anonymous):

do I have to use the quotient rule to figure out the derivative of \[\sqrt{x/2}\]

OpenStudy (lgbasallote):

nope.. youse constant multiple rule \[\sqrt{\frac x2} \implies \frac{\sqrt x}{\sqrt 2}\] it's just the same as \[\frac 1{\sqrt 2} (\sqrt x)\]

OpenStudy (lgbasallote):

so your coefficient is 1/sqrt 2... just take the derivative of sqrt x

OpenStudy (lgbasallote):

then multiply to 1/sqrt 2

OpenStudy (lgbasallote):

anyway... i have to leave now...so i hope you luck

OpenStudy (anonymous):

ahhh...thank you!!!

OpenStudy (lgbasallote):

welcome